What is the value of x after each of these statements is encountered in a computer program, if x =3 before the statement is reached? 


1
Expert's answer
2021-04-29T16:56:34-0400

i) if x +2=5 then x := 3*x +5

Since "3+2=5" is a true statement, after the command "if x +2=5 then x := 3*x +5" the value of "x" is "3\cdot 3+5 =14."

Answer: "x = 14"


ii) if (x +1=4) OR (2x +2=3) then x := x +1

Since "3+1=4" is a true statement, the statement "(3 +1=4) \text{ OR }(2\cdot 3 +2=3)" is also true, and thus after the command "if (x +1=4) OR (2x +2=3) then x := x +1" the value of "x" is "3+1=4"

Answer: "x = 4"